At the beginning of an experiment, a culture contains 1000 bacteria . Five hours later, there are 7600 bacteria. Assuming that the bacteria grow exponentially, how many bacteria will there be after 24 hours?
Answer by nerdybill(7384)

At the beginning of an experiment, a culture contains 1000 bacteria . Five hours later, there are 7600 bacteria. Assuming that the bacteria grow exponentially, how many bacteria will there be after 24 hours?
.
Exponential growth equation:
p = e^(kt)
where
p is amount after time t
k is a constant
t is time
.
From:"At the beginning of an experiment, a culture contains 1000 bacteria . Five hours later, there are 7600 bacteria." we get:
7600 = 1000e^(5k)
7600/1000 = e^(5k)
7.600 = e^(5k)
ln(7.600) = 5k
ln(7.600)/5 = k
0.40563 = k
.
Our "general equation" for growth then is:
p = e^(0.40563t)
which we use to solve:
how many bacteria will there be after 24 hours?
p = e^(0.40563*24)
p = 16900 (answer)